find the value of sin x if x is in the standard position and the point (-3,2) lies on its terminal side
Answer by lwsshak3(11628)

find the value of sin x if x is in the standard position and the point (-3,2) lies on its terminal side
This places the reference angle x in quadrant II where sinx=2/√(2^2+3^2)=2/√13